### v2

# Self-induced marginality in plastically deformed crystals

Abstract
Quasi-brittle plastic yielding is a salient feature of well-annealed glassy materials. Here we show that the same behavior is characteristic of perfect crystals after they experience mechanically driven elastic instability leading to massive nucleation of dislocations. We argue that such ’preparation’ effectively converts an atomic configuration from crystalline to quasi-amorphous. To understand the nature of the subsequent intermittent mechanical response we study a model 2D crystal subjected to AQS driving and show that both pre- and post-yield dislocation avalanches exhibit power law statistics with similar exponents indicative of self-induced marginal stability.
The complexity of plastic response of strain-driven crystalline solids is revealed by intermittent stress fluctuations accompanied by scale-free spatial organization of dislocations  Papanikolaou et al. (2017); Hochrainer et al. (2014); Derlet and Maaß (2016); Ispánovity et al. (2010). The same general features are also typical of amorphous plasticity  Berthier et al. (2024); Shang et al. (2020); Singh et al. (2020); Lin et al. (2014). To build a conceptual link between the two phenomena, we study in this Letter a relatively transparent prototypical example of a 2D square crystal subjected to athermal quasistatic (AQS) proportional loading by shear strain applied along one of the slip directions. To prepare a ’generic’ crystal we start from a pristine dislocation-free crystal and shear it it to the point of mechanical instability which is resolved through massive dislocation nucleation. We argue that such ’preparation’ effectively converts an atomic configuration from crystalline to quasi-amorphous, e.g. Ovaska et al. (2017). This term is used to refer to a solid that is structurally crystalline but mechanically amorphous in the sense that its mechanical response is dominated by non-affine relaxations, e.g. Zaccone and Scossa-Romano (2011).
We observe that plastic flow of such quasi-amorphous crystals begins with a hardening stage involving small spatially localized dislocation rearrangements. This ’microplasticity’ regime is terminated abruptly with a quasi-brittle yield leading to the formation of multi-grain pattern dominated by global shear bands. The post-yield plastic flow is characterized by a stress plateau with superimposed fluctuations representing broadly distributed dislocation avalanches. Similar quasi-brittle yield has been observed in well-annealed glasses Tanguy (2021); Parley and Sollich (2024); Ozawa et al. (2018). Despite the absence of dislocations in amorphous plasticity, both phenomena can be viewed as representing the evolution of elastic incompatibility with underlying elementary mechanical events represented by shear eigenstrains Picard et al. (2004); Budrikis et al. (2017); Baggioli et al. (2022).
Our main tool is a novel mesoscopic tensorial model (MTM) of crystal plasticity representing a conceptual trade-off between continuum and atomic descriptions Salman and Truskinovsky (2011, 2012); Baggio et al. (2019); Salman et al. (2021); Perchikov and Truskinovsky (2024); Baggio et al. (2023a, b); Zhang et al. (2020). Crucially, it allows one to capture the intermittent nature of dislocation avalanches while resolving topological (connectivity) restructuring involved in nucleation and annihilation of dislocations without any specialized phenomenological assumptions Weiss et al. (2021). In contrast to other computational approaches to micro-scale plasticity, e.g. Zepeda-Ruiz et al. (2021); Chan et al. (2010); Salvalaglio et al. (2020); Javanbakht and Levitas (2016); Beyerlein and Hunter (2016); Koslowski et al. (2002); Wang et al. (2003); Devincre et al. (2008); Ispánovity et al. (2014); Gómez-García et al. (2006); Po et al. (2014); Starkey et al. (2020); Zhang et al. (2025), the MTM allows one to deal with statistically meaningful number of dislocations while accounting in a geometrically exact way for finite deformations  Engel (1986); Pitteri and Zanzotto (2002).
The main idea of the MTM approach is to associate with material points an effective energy landscape whose tensorial periodicity accounts for crystallographically-invariant deformations. The resulting Landau-type model is characterized by infinitely many equivalent energy wells Ericksen (1970); Parry (1998); Folkins (1991); Conti and Zanzotto (2004); Bhattacharya et al. (2004); Salman and Truskinovsky (2011); Baggio et al. (2019); Perchikov and Truskinovsky (2024) and plastically deformed crystals emerge as a multi-phase mixtures with dislocations playing the role of effective domain boundaries. The small scale regularization is achieved by mesoscopic discretization. Despite operating with engineering concepts of stress and strain the resulting computational code correctly captures not only long-range elastic interactions of dislocations but also describes adequately their short-range interactions involved in the formation of complex dislocational entanglements, see also variations on the same theme in Denoual et al. (2010); Gao et al. (2019); Biscari et al. (2015); Arbib et al. (2023). We also note that the 2D version of the MTM, used in our numerical experiments, has some formal similarities with a single-slip-plane phase field model of Peierls-Nabarro type, which was found fully adequate to represent the critical nature of plastic flows, e.g. Koslowski et al. (2002, 2004). The main difference of our approach is the use of finite strain bulk elasticity compatible with multi-slip and capturing adequately lattice rotations which produced different values of critical exponents revealing the marginal stability of the underlying configurations.
Refer to caption Refer to caption
(a) (b)
Figure 1: (a) $GL(2,\mathbb{Z})$-induced tessellation of the configuration space of metric tensors $C_{11},C_{22},C_{12}$ with $\det\mathbf{C}=1$ stereographically projected on the Poincaré disk $p^{2}+q^{2}<1$ where $p=t(C_{11}-C_{22})/2$, $q=tC_{12}$, and $t=2/(2+C_{11}+C_{22})$. The highlighted domain $\mathscr{D}$ is the fundamental (minimal periodicity) domain. (b) Representative energy landscape with colors indicating the energy level. Only the lowest energy levels are visible with the bottoms of the wells corresponding to equivalent replicas of the reference square lattice.
Consider the continuum deformation field $\bf y=\bf y(\bf x),$ where $\bf y$ and $\bf x$ are positions of material points in the current and reference configurations, respectively. Due to frame indifference constraint, the strain energy density $\varphi$ can depend on the deformation gradient $\bf F=\nabla\bf y$ only through the metric tensor ${\bf C}={\bf F}^{T}{\bf F}$ which then plays the role of a tensorial order parameter. In MTM we assume that the function $\varphi(\bf C)$ additionally respects the symmetry of an underlying Bravais lattice. Such symmetry is assumed to be global extending beyond the conventional point group and capturing lattice invariant shears. In 2D it reduces to the constraint that $\varphi({\bf C})=\varphi({\bf m}^{T}{\bf C}{\bf m}),$ where unimodular integer valued matrices $\bf m$ belong to $GL(2,\mathbb{Z})$ SM . In view of this symmetry, the surface $\det{\bf C}=1$ in the 3D space $(C_{11},C_{22},C_{12})$ is tessellated into an infinite number of equivalent periodicity domains; if the energy density $\varphi({\bf C})$ is known in one of such domains, it can be extended to the whole configurational space of tensors $\bf C$ by the symmetry. To visualize the implied tessellation, it is convenient to project stereographically the infinite surface $\det{\bf C}=1$ on a unit (Poincaré) disk, see Fig. 1(a). In our numerical experiments we used the simplest piece-wise polynomial function $\varphi(\bf C)$ proposed in Conti and Zanzotto (2004), see SM for details. In Fig. 1(b) we illustrate the corresponding multi-well energy landscape. Here the equivalent energy wells, which are somewhat distorted due to strereographic projection, represent the equivalent replicas of the reference square lattice. A characteristic feature of such a landscape is a network of low-energy valleys corresponding in our case to simple shears parallel to crystallographic slip planes. These valleys represent conventional plastic ’mechanisms’ and are assumed to be flat in the classical continuum crystal plasticity.
Since the MTM preserves the energy barriers separating individual wells along the low-energy valleys, the function $\varphi(\mathbf{C})$ ends up being highly nonconvex. The corresponding continuum scale-free Landau model is then degenerate and needs to be regularized through the introduction of a cut-off length scale. This is achieved through the projection of the continuum problem on a uniform mesoscale grid while associating the nonconvex elastic energy to finite discrete elements. In other words, the infinite dimensional space of continuum deformation fields is reduced to a finite dimensional set of compatible, piece-wise affine mappings, with the linear scale of an individual finite element becoming a physical parameter SM . An incremental elastic energy minimization in a system of this type subjected to quasistatic loading will generically lead to a rich repertoire of instabilities with jump discontinuities representing elastic branch switching events. The energy losses associated with such jumps contribute to plastic dissipation; in the continuum limit the jumps will merge giving rise to rate-independent plasticity Puglisi and Truskinovsky (2005).
In our numerical experiments a square crystal was represented by $N\times N$ nodes, with $N=100,200,400$. The nodes were arranged into triangular finite elements. We loaded the system quasi-statically by applying the affine displacement field ${\bf u}(\alpha,{\bf x})=(\bar{\bf F}(\alpha)-\mathbb{1}){\bf x},$ with $\bar{\bf F}(\alpha)={\bf I}+\alpha{\bf e}_{1}\otimes{\bf e}_{2},$ where ${\bf e}_{i}$, $i=1,2$, is the orthonormal basis of the reference square lattice. In this way we imposed simple shear along one of the principal slip directions with shear amplitude $\alpha$ serving as the loading parameter. The loading was advance in strain increments of order $10^{-6}$ and after each increment the displacement field was updated by the energy minimization algorithm. Details of the discretization, of the incremental energy minimization algorithm and of the implementation of the loading protocol under periodic boundary conditions can be found in SM .
Figure 2: Dislocated configuration emerging after a homogeneous lattice is mechanically driven to the threshold of elastic instability. Colors indicate the level of the $\sigma_{xy}$ component of the Cauchy stress tensor. The highest stress levels of both signs correspond to the location of dislocation cores. Here $N=400$.
To obtain a ’generic’ sample, we first loaded a pristine (defect-free) crystal till the point of elastic instability of the affine configuration. The recorded critical value of the loading parameter $\alpha=\alpha_{c}\approx 0.138$ was found to be very close to the threshold of positive definiteness of the acoustic tensor Grabovsky and Truskinovsky (2013), see the explicit formulas in SM . To trigger the a generic instability we imposed a small Gaussian quenched disorder which was subsequently removed. The breakdown of an elastic state took the form of a massive nucleation of dislocations resulting in a highly inhomogeneous configuration of the type shown in Fig. 2. It is natural to assume that the unfolding of the elastic instability was arrested at a state which is only marginally stable and is therefore far from being random.
Figure 3: (a) Stress-strain responses of four different ’generic’ quasi-amorphous crystals of the type shown in Fig. 2; solid black curve indicates the average response. The stress-strain curves do not start from the origin because the global instability does not relaxes stresses to zero due to the emergence of self-equilibrated (residual) dislocation configuration. (b) Dislocated configuration in a typical pre-yield state characterizing the stage of ’microplasticity’. (c,d) Dislocated configuration characterizing different stages of unfolding of the quasi-brittle system size event. Colors indicate the level of the $\sigma_{xy}$ component of the Cauchy stress tensor. In (a-d) N=400.
To study the mechanical response of the obtained ’generic’ sample we continued to load it quasistatically while recording the ensuing stress-strain relation. The results are shown in Fig. 3(a) for four different realization of disorder with the solid black curve indicating the average response. Its salient feature is the presence of irregularly placed elastic branches interrupted by intermittent stress drops; the latter represent plastic avalanches involving partial unlocking of dislocation structures with superimposed distributed nucleation/anihilation of dislocations. A major diffuse stress drop, which we interpret as the indication of the beginning of plastic yield, separates two markedly different regimes. While the quasi-elastic pre-yield regime is characterized by finite overall rigidity, the post-yield regime is represented by an effective stress plateau; in both regimes we observe intense plastic fluctuations, see Fig. 3(a).
Our Fig. 3(b) illustrates the stress field in a typical pre-yield state $A$. Comparing to Fig.  2, the overall stress level is higher but the restructuring of the dislocation pattern is relatively minor with dislocations mostly displaced between the preexisting locking sites. Such ’microplasticity’ response is a characteristic pre-yield feature for both amorphous and (defective) crystalline solids Maaß and Derlet (2018); Papanikolaou et al. (2017); Sparks et al. (2019); Rizzardi et al. (2022) This regime ends with the system-size event which takes the form of single quasi-brittle plastic stress drop in well-annealed glasses and some sub-micron crystals Cui et al. (2017); Greer and Hosson (2011); Sharma et al. (2018); Mordehai et al. (2018). In Fig. 3(c,d) we show the snapshots of the spatial stress configurations in the intermediate state $B$ and the final state $C$ ending the transition to post-yield regime. The comparison of Fig. 3(b-d) shows that the associated global restructuring involves collective dislocation activity leading to the formation of system-spanning shear bands.
Additional aspects of the state $C$ can be seen in Fig. 4 where instead of the stress field, as in our Fig. 3(d), we now show the strain-energy density field. In this representation the locations of individual dislocations become visible. The general pattern is the development of low-energy patches of the original lattice forming polycrystalline grain texture with elastic energy localized around dislocation-rich grain boundaries. In the inset presented in Fig. 4 we show the zoom in on the configuration of elastic elements; individual dislocations were identified via Delaunay triangulation and the nodes with five/seven neighbors were shown in blue/red. In particular, we see the ubiquitous presence of $\Sigma 5$ grain boundaries which can be explained by the fact that the misorientation of neighboring grains is controlled by the overall compatibility of the deformation field Baggio et al. (2023b).
Figure 4: The micro-configuration configuration corresponding to the state $C$ marked in Fig. 3. Colors indicate the magnitude of strain-energy density. The inset shows a single grain and details the dislocation structure of the grain boundaries. Here N=400.
The ensuing post-yield regime, characterized by the relative stabilization of the average stress level, can be broadly interpreted as a quasi-stationary non-equilibrium steady state. A representative snapshot of the micro-configuration of the crystal in the state $D$ is shown in Fig. 5(a) where we see progressive maturation of the polycrystalline grain texture. During intermittent avalanches characterizing this regime larger grains occasionally merge while smaller grains continue to emerge. In this way the self-organization of dislocations continues to generate additional scales which contribute to the formation of the global hierarchical structure. The complexity of the emerging configuration is illustrated in Fig. 5(b) where we show the corresponding strain distribution in the Poincaré disk. While in the reference state (pristine crystal) all configurational points were concentrated in the origin, now they are spread all over the configurational space with expected accumulation near the bottoms of at least five energy wells representing different replicas of the unstressed square lattice; for comparison, in the pre-yield regime the system explores at most three energy wells immediately adjacent to the reference state. In other words, plastic yield not only marks the transitions from isolated dislocation motion to collective dislocation behavior, but also implies the access to a much broader repertoire of relaxation mechanisms. For instance, a more detailed analysis of the avalanche structure illustrated in SM shows that while in the pre-yield regime plastification takes place primarily in the form of isolated, linear arrangements of transformed elements, a typical post-yield avalanche reveals extended plastified regions with complex branching.
Refer to caption Refer to caption
(a) (b)
Figure 5: (a) Shear band development in the regime of stationary post-yield plastic flow represented by the state $D$ in Fig. 3. Here N=400. Colors indicate the level of the shear component of the Cauchy stress tensor $\sigma_{xy}$. (b) Corresponding distribution of local values of the metric tensors in the configurational space projected on the Poincaré disk. Background color represents the energy landscape whose 3D representation was shown in Fig. 1(b). The color of the gray dots represents the number of the finite elements sharing the same metric tensor.
An important window into the micro-mechanics of solids undergoing plastic flow is provided by the study of the statistical structure of plastic fluctuations Papanikolaou et al. (2017); Weiss et al. (2015); Alava et al. (2014). In our case the relevant observables would be the energy dissipated at a single avalanche $\Delta W$ and the corresponding macroscopic stress drop $\Delta\Sigma=A\Delta\sigma_{xy}$, where $A$ is the surface area of the sample. It is known that in amorphous solids the probability distribution of these observables is compatible with the finite-size scaling ansatz $P(s;L)\;=\;s^{-x}\,\mathcal{F}\!\left(\frac{s}{L^{D_{x}}}\right),$ where $L=Nh$ is the linear system size, $h$ is the size of an element, $x$ is the power-law exponent with $x=\tau$ for stress drops and $x=\epsilon$ for dissipated energy and $D_{x}$ is the corresponding fractal dimension ($D_{\epsilon}$ or $D_{\tau}$, respectively). Finally $\mathcal{F}(u)$ is a universal scaling function. It is also known that in amorphous plasticity the exponents $\epsilon$ and $\tau$ are similar for pre-yield and post-yield regimes while the exponents $D_{\epsilon}$ and $D_{\tau}$ may be different Tyukodi et al. (2016); Budrikis et al. (2017); Lerner et al. (2018); Ozawa et al. (2022).
To determine these power-law exponents for our numerical experiments we employed a robust data-collapse optimization procedure detailed in SM . The resulting data collapse is illustrated in Fig.  6. One can see that for our ’generic’ quasi-amorphous crystals the statistics of avalanches follows the same pattern as for their amorphous counterparts. In particular, similar to the case of well-annealed glasses Papanikolaou et al. (2017); Weiss et al. (2021); Sethna et al. (2001); Ispánovity et al. (2014); Alava et al. (2014); Lehtinen et al. (2016), the scaling behavior remains basically unchanged across the quasi-brittle yielding transition.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 6: Optimal data collapses of the probability distributions onto universal master curves. (a, b) Collapse of the energy drops $\Delta W$ in the pre-yield generating exponents $\epsilon=0.9195$ and $D_{\epsilon}=0.3640$ (a) and post-yield regimes generating exponents $\epsilon=1.0207$ and $D_{\epsilon}=0.9836$ (b). Collapse of the macroscopic stress drops $\Delta\Sigma$ in the pre-yield generating exponents $\tau=0.9070$ and $D_{\tau}=0.3987$ (c) and post-yield regimes, generating exponents $\tau=1.2476$ and $D_{\tau}=1.4379$ (d).
More specifically, the obtained power law range for energy drops extends over up to six decades with pre-yield and post-yield energy distributions sharing almost the same value of the exponent $\epsilon=1.0$. The invariance of this exponent across the yielding transition suggests that while the spatial organization of plastic events changes drastically from local and non-cooperative to global and collective, the statistical distribution of energy drops remains insensitive to these details responding only to the global organization of the underlying energy landscape Shang et al. (2020); Oyama et al. (2021).
The value of the power-law exponent $\epsilon\approx 1.0$ is a signature of an archetypically ’wild’ crystal plasticity in the sense of Weiss et al. (2015). The same value was obtained in discrete dislocation dynamics studies of disorder-free 2D samples containing a fixed number of preexisting dislocations Ispánovity et al. (2014); Salmenjoki et al. (2021) and was also recorded in numerical experiments on pristine crystals using a scalar version of the MTM Zhang et al. (2020). The exponent $\epsilon\approx 1.0$ has been previously interpreted in crystal plasticity framework as representing either dislocation jamming or self-induced glassiness Ovaska et al. (2015); Lehtinen et al. (2016); Zhang et al. (2016).
It is rather striking that exactly the same value of the exponent has been also recorded in the studies of plastic flows in structural glasses Tyukodi et al. (2019); Ferrero and Jagla (2019); Shang et al. (2020). A theoretical understanding of the relation between the emergence of the power law exponent $\epsilon\approx 1.0$ and the structure of the underlying energy landscapes has been developed in the theory of spin glasses. Thus, it was shown that the hierarchical (ultrametric) organization of energy wells in the phase space results in an intermittent, scale-free response to quasistatic deformation  Franz and Spigler (2017); Müller and Wyart (2015); Berthier et al. (2019); Dennis and Corwin (2020). Moreover, it was rigorously proved that in the mean-field limit the associated distribution of energy avalanches must be of a power law with exponent $\epsilon\approx 1.0$. In the spin glass context it has been also understood that the reason behind this particular value of the exponent is the marginal stability of the system  Franz and Spigler (2017); Pázmándi et al. (1999); Doussal et al. (2012).
Since glasses can be viewed as liquids at a quenching threshold, their marginal stability is similar to marginality exhibited by granular matter at a jamming point. If such systems are mechanically driven, the proximity to unstable modes leads to the mixing of statics (stability) and dynamics (instability) making the mechanical response inherently intermittent  Lamp et al. (2022); Rossi and Tarjus (2022). Similarly, since our ’quasi-amorphous crystal’ emerged from arrested dynamics, one can argue that the associated self-generated disorder brought the system from a solid to an effectively glassy state. Furthermore, our results suggest that the implied marginality is not affected by the quasi-brittle yield, which apparently does not compromise the global organization of the energy landscape while, of course, affecting the reachability of its different subdomains Regev et al. (2021).
The situation is not exactly the same for the exponent $\tau$ characterizing the power law statistics of the stress drops $\Delta\Sigma$. Note first that in the pre-yield regime, the power-law exponents for stress and energy are rather similar, $\tau\approx\epsilon$, which suggests that the dissipated energies and the stress drops are roughly proportional, $\Delta W\sim\Delta\Sigma$. Such linear scaling implies some kind of caging, which means a highly constrained regime of dislocation motion with individual avalanches arrested by a regular arrays of obstacles. This may be also the reason behind the strong hardening behavior in the pre-yield regime.
Instead, in the post-yield regime the recovered exponents $\tau$ and $\epsilon$ are markedly different. Here individual avalanches are no longer hindered by local structural constraints and one can assume that the stress drop is only constrained by the corresponding strain increment due to the constraints of (nonlinear) elasticity. Then, if we assume that $\Delta W\sim\Delta\Sigma^{\gamma}$ we can expect that $\epsilon-1=(\tau-1)/\gamma$. In particular, if elasticity is linear, $\gamma=2$, and the measured value is $\tau\approx 1.2$, we can expect that $\epsilon\approx 1.10$ which is in reasonable agreement with our measured value $\epsilon\approx 1.02$ given that the uncertainty for our relatively small system is of the order $\sim\pm 0.2$, see SM for details.
Note next that the pre-yield values of the cutoff parameters $L^{D_{x}}$ are much smaller than their post-yield values which indicates that larger avalanches become more probable  Lin et al. (2014); Ruscher and Rottler (2021). The emergent collective behavior, reflected in such broader extent of avalanche activity, reaches in post yield regime the size of the system suggesting the divergence of the characteristic length. Further differences between pre-yield and post-yield mechanical responses emerge from the analysis of the values of the exponents $D_{x}$ providing information about the morphology of the plastically deforming regions. Thus, in In the pre-yield regime the energy avalanches are characterized by a low fractal dimension $D_{\epsilon}\approx 0.36$, pointing towards highly localized, spatially scattered, macroscopically isolated rearrangements. Instead, in the post-yield regime we obtain the value $D_{\epsilon}\approx 0.98$ which suggests the emergence of 1D-like system spanning bands. The latter reflect highly cooperative nature of the underlying plastic deformation and the physical nature of the implied criticality is presently actively debated Grinstein (1995); Zippelius et al. (2023); Oyama et al. (2023); Xing et al. (2024); Zhang et al. (2020).
Finally we mention that the observed similarity between the structure of intermittent plastic fluctuations in amorphous and crystalline solids is supported by the fact that elasto-plastic models, typically used to model amorphous plasticity Nicolas et al. (2018); Rossi et al. (2022); Ferrero and Jagla (2019); Fernández Castellanos et al. (2021), are very similar to the MTM which operates within basically the same finite element setting. The difference is that the phenomenological yield thresholds of elasto-plastic models are replaced in the MTM by elastic instabilities originating from the non-convexity of the energy landscape. Accordingly, the fixed linear elastic propagators are replaced by the solution ’on the fly’ of the corresponding nonlinear elasticity problems. Note also that the crucial nonlinearity in the MTM is of universal geometrical nature as it originates from the exact description of finite elastic deformations. However, the observed general agreement in the values of the computed exponents suggests that we are dealing with the same universality class.
To conclude, we showed that mechanically driven perfect crystals can exhibit quasi-brittle plastic yielding which is remarkably similar to the behavior of well-annealed glassy materials. The implied parallel features of the mechanical response emerge after pristine crystals acquire self-generated disorder resulting from massive dislocation nucleation during the breakdown of an affine elastic configuration. The ensuing quasi-amorphous crystals exhibit pre- and post-yield avalanches with power law statistics whose matching exponents are indicative of the marginality of the associated mechanical system. Adaptation of the same model to realistic 3D crystals will allow one to distinguish between edge and screw dislocations opening the way to the adequate account of such physically important effects as dislocation climb, cross slip and forest hardening. To be adequate such an analysis should be, of course, performed with the account of the effects of finite temperature Zhang et al. (2023); Li et al. (2024).
Acknowledgments. The authors thank N. Gorbushin for helpful comments. The work of O.U.S. was supported by the grants ANR–21-CE08-MESOCRYSP, ANR-20-CE91-0010, ANR-24-CE91-0002. L. T. was supported by the grants ANR–21-CE08-MESOCRYSP and ERC-H2020-MSCA-RISE-2020-101008140.